OpenStudy (anonymous):

What is the surface area of the prism? A. 204 in2 B. 120 in2 C. 180 in2 D. 192 in2

OpenStudy (anonymous):

@LegendarySadist

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

what is the formula?

OpenStudy (anonymous):

rectangular prism formula= 2*b+p*h

OpenStudy (anonymous):

Find the area of the triangle and multiply it by 2. Then find the perimeter of the triangle and multiply it by 10

OpenStudy (anonymous):

area of the base......

OpenStudy (anonymous):

so multiply 10*8= 80 then multiply it by 2

OpenStudy (anonymous):

No, we're using the triangle as the base since it's much easier.

OpenStudy (anonymous):

160

OpenStudy (anonymous):

Chiko, you're doing it the hard way you know.

OpenStudy (anonymous):

im just doing it by the way they taught me...

OpenStudy (anonymous):

Well there's an easier way. Don't use the rectangle as the base. Use the triangle as the base. So \[(\triangle area*2)+(\triangle perimeter * 10)\] Find the triangle area and perimeter then plug it in for the answer.

OpenStudy (anonymous):

CONFUSED IM ABOUT TO DIE!

OpenStudy (anonymous):

so take the p and multiply it by 3 and and then add what you got to 160

OpenStudy (anonymous):

*Mind gets mixed up* wait what r we doing? IDK WHAT TO DO I FORGOT

OpenStudy (anonymous):

Do you know how to solve for the area of a triangle?

OpenStudy (anonymous):

No???Yes???IDK

OpenStudy (anonymous):

*starts to cry fr*

OpenStudy (anonymous):

160+ 18*3= your answer

OpenStudy (anonymous):

214!

OpenStudy (anonymous):

Angel, do you know how to solve for the area of a triangle? The area of a triangle is .5*height*base. So it will be \[.5*3*8\]

OpenStudy (anonymous):

not 1 of my answers

OpenStudy (anonymous):

4

OpenStudy (anonymous):

And chiko, that would get you 208, which is incorrect.

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

Ok angel, let's start from the beginning. let's find the area of the triangle, which will be\[\frac{ 1 }{ 2 }*(3*8)\]

OpenStudy (anonymous):

Sorry had to go outside and get some air. 12

OpenStudy (anonymous):

12 is the answer to that probem^

OpenStudy (anonymous):

problem*

OpenStudy (anonymous):

Please help! i have to go soon and i have 1 more question after this!

OpenStudy (anonymous):

Bye i gtg eat lunch ;(